Please help me. 400 people were at a baseball game. Student tickets are $2. Adult tickets are $3. Total ticket sales were $1050. How many of each tickets were sold.
Answer by srw491(2)

150 student tickets, 250 adult tickets
let x=student tickets, y=adult tickets
need 2 equations:
1) 2x+3y=1050
2) x+y=400
substitute: y=400-x into first equation
2x+3(400-x)=1050
2x+1200-3x=1050
-x=-150 => x=150 substitute back into second equation
150+y=400
y=250